OpenStudy (anonymous):

xy=4 a. no b. yes, xy=4 c. yes, y=4x d. yes, y-4x=0

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

y and x mean 1 but i get lost after that

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

Is this a problem from a textbook?

OpenStudy (anonymous):

it is a exam question out of my exam book

OpenStudy (anonymous):

Can you rewrite the problem word for word? Which course is this a problem from?

OpenStudy (anonymous):

algebra 1

OpenStudy (anonymous):

xy = 4 is the question and the answers are a. no b. yes , xy = 4 c. yes , y = 4x d. yes , y - 4x =0

OpenStudy (anonymous):

Is he looking for all the answers that are equivalent to the first equation? I'm going to assume that's what he wants. If x*y=4, then to move x to the other side of the equation, you divide both sides by x. Then you have y=4/x, so y=4x is not equivalent. If you want to move the 4 to the other side of the equation, you subtract it from both sides getting (x*y)-4=0, so y-4x is not equivalent. x*y=4, is equivalent to x*y=4, assuming x and y remain the same variables. I hope this helps, I'm really not sure what else he could be asking

OpenStudy (anonymous):

so are you saying the answer is B and thank you cause someone else was explaining it the way you just did so i think your right im just not good in math.

OpenStudy (anonymous):

The wording is a bit confusing. If that is all the information your teacher gave you, I'm inclined to think that your troubles may be the fault of your instructor. Don't put yourself down, look for someone at your school who can help you one-on-one. Believe me, an instructor that you click with is more important to learning than you can imagine.

OpenStudy (anonymous):

it says for the problem you just helped me with( determine if the equation is linear and if it is then put it in standard form)

OpenStudy (anonymous):

Okay! That helps a lot. Now I understand what it is asking.

OpenStudy (anonymous):

thanks for helping me

OpenStudy (anonymous):

So lets start by understanding what a linear equation is. What is your best guess?

OpenStudy (anonymous):

to be honest i dont know i dropped out of school in the middle of learning this stuff and now since im back doing homeschool im pretty much lost i use to be really good in math

OpenStudy (anonymous):

Well, what does the word linear look like? It contains the word "line", which is important because a line is exactly what a linear equation looks like!

OpenStudy (anonymous):

So, a cartesian coordinate system, is a system for graphing mathematical equations. http://www.mathsisfun.com/data/cartesian-coordinates.html On a cartesian coordinate system, a linear equation will create a straight line.

OpenStudy (anonymous):

The way it does this, is by growing, or shrinking at a constant, or consistent rate. In other words, it starts at one point, such as x=0, y=0, picks a rate, such as 1 and continues at that rate forever. So if the rate its going is 1, it will go up one every time it goes over one. By "over one" I mean, moves to the right. So if it starts at x=0, y=0, this little point, let's call him Harry, will move over 1 to x=1 and jump up 1 to y=1. He will continue to do this forever. So next he moves over to x=2 and jumps up to y=2. Harry is moving in a linear pattern

OpenStudy (anonymous):

Do you follow me so far?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Okay, so linear equations are in the form of y=mx+b x is the independent variable. you pick whatever value* you want for x. y is the dependant variable. Its value depends on the x value. m is the rate at which Harry jumps up for every 1 x he moves over b is where Harry starts *Almost any value, but that's a different lesson

OpenStudy (anonymous):

Your teacher wants you to put x*y=4 into the form y=mx+b

OpenStudy (anonymous):

oh okay well thanks for helping me i think im starting to understand it better.

OpenStudy (anonymous):

Would you like me to continue?